# Hao Shi1  Bin Xie2  Yingfei Liu2  Lin Sun4  Fengrong Liu5  Tiancai Wang2 Erjin Zhou2  Haoqiang Fan2  Xiangyu Zhang3,6  Gao Huang1🖂

### v1

@@ 4.2 Simulated Evaluation on SimplerEnv
We evaluate on two SimplerEnv suites: Bridge and Fractal. For SimplerEnv-Bridge, we train on Bridge v2 dataset (Walke et al., 2023) for 50k steps with validation every 2.5k steps. Results are reported at the best validation step, and each task is evaluated with 24 trials to compute success rates. For SimplerEnv-Fractal, we train on the RT-1 dataset (Brohan et al., 2022) for 80k steps with validation every 5k steps. Evaluation covers Visual Matching (VM) and Visual Aggregation (VA) settings. VM mirrors the real setup to reduce sim-to-real gap, while VA stress-tests robustness by altering background, lighting, distractors, and table textures. The Fractal testbed includes 336 variants, yielding 2,352 trials in total.
@@
As shown in Tab. 1, MemoryVLA achieves an average success rate of 71.9%, a +14.6 point gain over the CogACT-Large baseline, and surpasses recent state-of-the-art VLAs including $\pi_{0}$ (Black et al., 2024). Per task, success rates are 75.0% on Spoon on Towel, 75.0% on Carrot on Plate, 37.5% on Stack Cube, and 100.0% on Eggplant in Basket.
@@
Table 3: Performance comparison on LIBERO (Liu et al., 2023a) with Franka robot. Success rates (%) are reported across five suites. * indicates methods using additional proprioceptive and wrist-camera inputs. CogACT results are reproduced by us. For methods without LIBERO-90 results, we report the average over the first four suites.
Method Spatial Object Goal Long LIBERO-90 Avg. Success
Diffusion Policy (Chi et al., 2023) 78.3 92.5 68.3 50.5 72.4
Octo (Team et al., 2024) 78.9 85.7 84.6 51.1 75.1
MDT (Reuss et al., 2024) 78.5 87.5 73.5 64.8 76.1
UniACT (Zheng et al., 2025b) 77.0 87.0 77.0 70.0 73.0 76.8
MaIL (Jia et al., 2024) 74.3 90.1 81.8 78.6 83.5
SpatialVLA (Qu et al., 2025) 88.2 89.9 78.6 55.5 46.2 71.7
TraceVLA (Zheng et al., 2024b) 84.6 85.2 75.1 54.1 74.8
OpenVLA (Kim et al., 2024) 84.7 88.4 79.2 53.7 73.5 75.9
CoT-VLA (Zhao et al., 2025) 87.5 91.6 87.6 69.0 81.1
$\pi_{0}$-FAST* (Pertsch et al., 2025) 96.4 96.8 88.6 60.2 83.1 85.0
TriVLA (Liu et al., 2025c) 91.2 93.8 89.8 73.2 87.0
4D-VLA (Zhang et al., 2025a) 88.9 95.2 90.9 79.1 88.6
CogACT (Li et al., 2024a) 97.2 98.0 90.2 88.8 92.1 93.2
$\pi_{0}$(Black et al., 2024) 96.8 98.8 95.8 85.2 94.2
MemoryVLA (Ours) 98.4 98.4 96.4 93.4 95.6 96.5(+3.3)
@@ 4.3 Simulated Evaluation on LIBERO
We evaluate on the LIBERO (Liu et al., 2023a) benchmark with a Franka robot across five suites: Spatial, Object, Goal, Long, and LIBERO-90. The first four suites contain 10 tasks each, and LIBERO-90 contains 90 tasks. Following OpenVLA (Kim et al., 2024), 50 demonstrations per task are used. Separate models are trained for Spatial, Object, and Goal for 20k steps each, while Long and LIBERO-90 are trained jointly for 40k steps. Validation is performed every 1k steps and results are reported at the best validation step. Each task is evaluated with 50 trials and per-suite average success rates are reported.
@@
As shown in Tab. 3, MemoryVLA achieves an overall success rate of 96.5%, improving CogACT by +3.3 points and surpassing $\pi_{0}$. Per-suite success rates are 98.4% on Spatial, 98.4% on Object, 96.4% on Goal, 93.4% on Long, and 95.6% on LIBERO-90. Note that MemoryVLA uses only third-person RGB, without wrist views or proprioceptive states.
@@
Method
General Tasks
Insert
Circle
Egg
in Pan
Egg
in Oven
Stack
Cups
Stack
Blocks
Pick Diverse
Fruits
Avg.
Success
OpenVLA (Kim et al., 2024) 47 27 53 40 13 4 31
$\pi_{0}$ (Black et al., 2024) 67 73 73 87 53 80 72
CogACT (Li et al., 2024a) 80 67 60 93 80 76 76
@@
Method
Long-horizon Temporal Tasks
Seq. Push
Buttons
Change
Food
Guess
Where
Clean Table
& Count
Pick Place
Order
Clean Rest.
Table
Avg.
Success
OpenVLA (Kim et al., 2024) 6 3 0 15 27 0 9
$\pi_{0}$ (Black et al., 2024) 25 42 24 61 82 80 52
CogACT (Li et al., 2024a) 15 47 40 67 90 84 57
@@
Variant
Avg.
Success
Memory
Type
Cognitive Mem. 63.5
@@ 4.4 Real-world Evaluation
As shown in Tab. 4, MemoryVLA achieves average success scores of 85% on the six General tasks and 83% on the six Long-horizon Temporal tasks, exceeding CogACT by +9 and +26 percentage points, respectively, and surpassing $\pi_{0}$ across both suites. On General tasks, we match or exceed the strongest baseline on every task, with notable gains on Egg in Pan (+13) and Egg in Oven (+20). On Long-horizon Temporal tasks, improvements are larger, including +43 on Seq. Push Buttons, +38 on Change Food, +32 on Guess Where, and +17 on Clean Table & Count. These results indicate strong real-world competence on general manipulation and highlight the benefits of temporal memory for long-horizon control.
@@ 5 Conclusion
[…]
$\etocsettocdepthsubsection \printcontents[app]$l0
Appendix A Robustness and Generalization Evaluation
A.1 Real-world Evaluation
@@ A.2 Simulation Evaluation
Appendix B Additional Training Details
B.1 Hyper-parameters
@@ B.2 Training Data
For the SimplerEnv-Bridge benchmark, we train on BridgeData v2 (Walke et al., 2023), a large-scale, language-conditioned real-robot manipulation dataset of roughly 60,000 teleoperated trajectories collected on WidowX robots across diverse tabletop settings. Episodes pair language instructions with demonstrations of skills such as picking, placing, pushing, stacking, and folding.
@@
For the SimplerEnv-Fractal benchmark, we use RT-1 (Brohan et al., 2022), a large-scale real-world dataset of roughly 130,000 episodes spanning 700+ tasks, collected over 17 months by the Google Robot fleet and paired with natural-language instructions.
LIBERO
LIBERO (Liu et al., 2023a) provides simulation tasks with a Franka robot across five suites: Spatial, Object, Goal, Long, and LIBERO-90, totaling 130 language-conditioned tasks. Each task supplies 50 demonstrations.
@@ B.3 Training Setup
Models are trained for 50k steps on Bridge v2, the dataloader is implemented as a streaming queue: each episode is pushed as a sequence of frames with an attached episode ID, and batches are constructed by popping consecutive samples. Within a batch, frames come from a single episode whenever possible; if an episode ends before the batch is filled, the remaining slots are completed with frames starting from the next episode. The memory length is fixed to 16.
@@
Following OpenVLA (Kim et al., 2024), we train with 50 demonstrations per task after removing failed trajectories from the dataset. Spatial, Object, and Goal suites are trained separately for 20k steps each, while Long-10 and LIBERO-90 (also referred to as Long-90) are treated as a single family of long-horizon data and trained jointly for 40k steps. The dataloader adopts a grouped sampling strategy: in each iteration, 16 frames are randomly sampled from within a single episode, matching the memory length of 16 used throughout training.
@@
Models are trained for 5k-20k steps depending on task and dataset size. The General suite contains 50-150 demonstrations per task, while Long-horizon Temporal tasks use 200-300 demonstrations. The dataloader follows the same design as in SimplerEnv-Bridge, but memory length differs by setting: 16 for General tasks and 256 for the longer-horizon Temporal tasks.
B.4 Data Augmentation
@@
Appendix C Additional Evaluation Details
C.1 SimplerEnv
Evaluation follows the official CogACT protocol (Li et al., 2024a). We adopt the same evaluation scripts and use the adaptive action ensemble strategy introduced in CogACT, with ensemble coefficient $\alpha=0.1$, ensemble horizon set to $7$ for Bridge and $2$ for Fractal. For Bridge, models are trained for 50k steps and validated every 2.5k steps, since the denoising objective of diffusion models does not reliably indicate policy quality, we report success rates at the best validation step. For Fractal, training runs for 80k steps with validation every 5k steps, and evaluation covers 336 variants in total (Tab. 10), we similarly report success rates at the best validation step, VM and VA settings are evaluated separately.
@@ C.1 SimplerEnv
C.2 LIBERO
Evaluation on LIBERO (Liu et al., 2023a) is conducted across all five suites (Spatial, Object, Goal, Long, and LIBERO-90). Models are validated every 1k steps, and each task is evaluated with 50 trials. Success rates are reported at the best validation step. Unlike SimplerEnv, no action ensemble strategy is used in our LIBERO experiments.

### v2

@@ 4.2 Simulated Evaluation on SimplerEnv
We evaluate on two SimplerEnv suites: Bridge and Fractal. For SimplerEnv-Bridge, we train on Bridge v2 dataset (walke2023bridgedata) for 50k steps with validation every 2.5k steps. Results are reported at the best validation step, and each task is evaluated with 24 trials to compute success rates. For SimplerEnv-Fractal, we train on the RT-1 dataset (brohan2022rt) for 80k steps with validation every 5k steps. Evaluation covers Visual Matching (VM) and Visual Aggregation (VA) settings. VM mirrors the real setup to reduce sim-to-real gap, while VA stress-tests robustness by altering background, lighting, distractors, and table textures. The Fractal testbed includes 336 variants, yielding 2,856 trials in total.
@@
As shown in Tab. 1, MemoryVLA achieves an average success rate of 71.9%, a +14.6 point gain over the CogACT-Large baseline, and surpasses recent state-of-the-art VLAs including $\pi_{0}$ (black2024pi_0). Per task, success rates are 75.0% on Spoon on Towel, 75.0% on Carrot on Plate, 37.5% on Stack Cube, and 100.0% on Eggplant in Basket.
@@
Table 1: Performance comparison on SimplerEnv-Bridge (li2024evaluating) with WidowX robot. CogACT-Large is our re-evaluated baseline using official weight, and MemoryVLA achieves a +14.6 gain in average success. Entries marked with * are reproduced from open-pi-zero, which leverage additional proprioceptive state inputs; they also adopt Uniform/Beta timestep sampling.
Method Spoon on Towel Carrot on Plate Stack Cube Eggplant in Basket Avg. Success
RT-1-X (o2024open) 0.0 4.2 0.0 0.0 1.1
OpenVLA (kim2024openvla) 4.2 0.0 0.0 12.5 4.2
Octo-Base (team2024octo) 15.8 12.5 0.0 41.7 17.5
TraceVLA (zheng2024tracevla) 12.5 16.6 16.6 65.0 27.7
RoboVLMs (liu2025towards) 45.8 20.8 4.2 79.2 37.5
SpatialVLA (qu2025spatialvla) 16.7 25.0 29.2 100.0 42.7
Magma (yang2025magma) 37.5 29.2 20.8 91.7 44.8
CogACT-Base (li2024cogact) 71.7 50.8 15.0 67.5 51.3
$\pi_{0}$-Uniform* (black2024pi_0) 63.3 58.8 21.3 79.2 55.7
CogACT-Large (li2024cogact) 58.3 45.8 29.2 95.8 57.3
CronusVLA (li2025cronusvla) 66.7 54.2 20.8 100.0 60.4
$\pi_{0}$-Beta* (black2024pi_0) 84.6 55.8 47.9 85.4 68.4
MemoryVLA (Ours) 75.0 75.0 37.5 100.0 71.9(+14.6)
Table 2: Performance comparison on SimplerEnv-Fractal (li2024evaluating) with Google robot. Success rates (%) are reported for Visual Matching (VM) and Visual Aggregation (VA) suites. MemoryVLA achieves an overall +4.6 gain over CogACT. O./C. denotes Open/Close, and * follow Tab. 1.
Method Visual Matching (VM) Visual Aggregation (VA) Overall
Coke Can Move Near O. / C. Drawer Put in Drawer Avg. Coke Can Move Near O. / C. Drawer Put in Drawer Avg.
Octo-Base (team2024octo) 17.0 4.2 22.7 0.0 11.0 0.6 3.1 1.1 0.0 1.2 6.1
RT-1-X (o2024open) 56.7 31.7 59.7 21.3 42.4 49.0 32.3 29.4 10.1 30.2 36.3
OpenVLA (kim2024openvla) 18.0 56.3 63.0 0.0 34.3 60.8 67.7 28.8 0.0 39.3 36.8
RoboVLMs (liu2025towards) 76.3 79.0 44.9 27.8 57.0 50.7 62.5 10.3 0.0 30.9 44.0
TraceVLA (zheng2024tracevla) 45.0 63.8 63.1 11.1 45.8 64.3 60.6 61.6 12.5 49.8 47.8
RT-2-X (o2024open) 78.7 77.9 25.0 3.7 46.3 82.3 79.2 35.5 20.6 54.4 50.4
Magma (yang2025magma) 75.0 53.0 58.9 8.3 48.8 68.6 78.5 59.0 24.0 57.5 53.2
SpatialVLA (qu2025spatialvla) 79.3 90.0 54.6 0.0 56.0 78.7 83.0 39.2 6.3 51.8 53.9
$\pi_{0}$-Uniform* (black2024pi_0) 88.0 80.3 56.0 52.2 69.1
$\pi_{0}$-Beta* (black2024pi_0) 97.9 78.7 62.3 46.6 71.4
CogACT (li2024cogact) 91.3 85.0 71.8 50.9 74.8 89.6 80.8 28.3 46.6 61.3 68.1
@@ 4.3 Simulated Evaluation on LIBERO
We evaluate on the LIBERO (liu2023libero) benchmark with a Franka robot across five suites: Spatial, Object, Goal, Long-10, and Long-90. The first four suites contain 10 tasks each, and Long-90 contains 90 tasks. Following OpenVLA (kim2024openvla), 50 demonstrations per task are used. Separate models are trained for Spatial, Object, and Goal for 20k steps each, while Long-10 and Long-90 are trained jointly for 40k steps. Validation is performed every 1k steps and results are reported at the best validation step. Each task is evaluated with 50 trials and per-suite average success rates are reported.
@@
As shown in Tab. 3, MemoryVLA achieves an overall success rate of 96.5%, improving CogACT by +3.3 points and surpassing $\pi_{0}$. Per-suite success rates are 98.4% on Spatial, 98.4% on Object, 96.4% on Goal, 93.4% on Long-10, and 95.6% on Long-90. Note that MemoryVLA uses only third-person RGB, without wrist views or proprioceptive states.
Table 3: Performance comparison on LIBERO (liu2023libero) with Franka robot. Success rates (%) are reported across five suites. * indicates methods using additional proprioceptive and wrist-camera inputs. CogACT results are reproduced by us. For methods without LIBERO-90 results, we report the average over the first four suites.
Method Spatial Object Goal Long-10 Long-90 Avg. Success
Diffusion Policy (chi2023diffusion) 78.3 92.5 68.3 50.5 72.4
Octo (team2024octo) 78.9 85.7 84.6 51.1 75.1
MDT (reuss2024multimodal) 78.5 87.5 73.5 64.8 76.1
UniACT (zheng2025universal) 77.0 87.0 77.0 70.0 73.0 76.8
MaIL (jia2024mail) 74.3 90.1 81.8 78.6 83.5
SpatialVLA (qu2025spatialvla) 88.2 89.9 78.6 55.5 46.2 71.7
TraceVLA (zheng2024tracevla) 84.6 85.2 75.1 54.1 74.8
OpenVLA (kim2024openvla) 84.7 88.4 79.2 53.7 73.5 75.9
CoT-VLA (zhao2025cot) 87.5 91.6 87.6 69.0 81.1
$\pi_{0}$-FAST* (pertsch2025fast) 96.4 96.8 88.6 60.2 83.1 85.0
CronusVLA (li2025cronusvla) 90.1 94.7 91.3 68.7 86.2
TriVLA (liu2025trivla) 91.2 93.8 89.8 73.2 87.0
4D-VLA (zhang20254d) 93.8 92.8 95.6 86.5 92.2
CogACT (li2024cogact) 97.2 98.0 90.2 88.8 92.1 93.2
$\pi_{0}$(black2024pi_0) 96.8 98.8 95.8 85.2 94.2
MemoryVLA (Ours) 98.4 98.4 96.4 93.4 95.6 96.5(+3.3)
Table 4: Performance comparison on Mikasa-Robo (cherepanov2025memory) with Franka robot. Success rates (%) are reported. CronusVLA results are reproduced by us.
Model ShellGame Intercept Remb. Remb. Remb. Avg.
Touch Medium Color3 Color5 Color9 Success
CronusVLA (li2025cronusvla) 32 5 31 13 9 18.0
SpatialVLA (qu2025spatialvla) 23 27 27 17 11 21.0
OpenVLA-OFT (kim2025fine) 47 14 59 16 6 28.4
PI-0 (black2024pi_0) 33 42 35 22 15 29.4
MemoryVLA (Ours) 88 24 44 30 20 41.2(+11.8)
4.4 Simulated Evaluation on Mikasa-Robo
Training and Evaluation Setup
We evaluate on the Mikasa-Robo (cherepanov2025memory) benchmark. In our experiments, we compare MemoryVLA with VLA-style baselines used in Mikasa-Robo, including OpenVLA-OFT (kim2025fine), PI-0 (black2024pi_0), and SpatialVLA (qu2025spatialvla). We additionally reproduce CronusVLA (li2025cronusvla), a contemporaneous VLA model that uses temporal context. Following the standard Mikasa-Robo protocol using 250 demonstrations per task at 128×128 resolution, 100 evaluation episodes per task and end-effector control. Standard 5 tasks are trained jointly for 20k steps, and validation is performed every 1k steps and results are reported at the best validation step.
Evaluation Results on Mikasa-Robo
MemoryVLA consistently achieves the highest performance across all tasks, with an average of 11.8% improvement over the previous state-of-the-art, especially +41.0% on ShellGameTouch task. The results are shown in Tab. 4.
@@ 4.4 Simulated Evaluation on Mikasa-Robo
Method General Tasks
Insert Circle Egg in Pan Egg in Oven Stack Cups Stack Blocks Pick Diverse Fruits Avg. Success
OpenVLA (kim2024openvla) 47 27 53 40 13 4 31
$\pi_{0}$ (black2024pi_0) 67 73 73 87 53 80 72
CogACT (li2024cogact) 80 67 60 93 80 76 76
@@
Method Long-horizon Temporal Tasks
Seq. Push Buttons Change Food Guess Where Clean Table & Count Pick Place Order Clean Rest. Table Avg. Success
OpenVLA (kim2024openvla) 6 3 0 15 27 0 9
$\pi_{0}$ (black2024pi_0) 25 42 24 61 82 80 52
CogACT (li2024cogact) 15 47 40 67 90 84 57
@@
Variant Avg. Success
Memory Type Cognitive Mem. 63.5
@@ 4.5 Real-world Evaluation
As shown in Tab. 5, MemoryVLA achieves average success scores of 85% on general tasks and 83% on long-horizon temporal tasks, exceeding CogACT by +9 and +26 percentage points, respectively, and surpassing $\pi_{0}$ across both suites. On general tasks, it exceeds the strongest baseline on every task, with notable gains on Egg in Pan (+13) and Egg in Oven (+20). On long-horizon temporal tasks, improvements are larger, including +43 on Seq. Push Buttons, +38 on Change Food, +32 on Guess Where, and +17 on Clean Table & Count. These results demonstrate strong real-world competence and highlight the benefits of temporal memory.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Science and Technology Major Project of China under Grant No. 2023ZD0121300, the National Natural Science Foundation of China under Grants U24B20173 and 62321005, the Scientific Research Innovation Capability Support Project for Young Faculty under Grant ZYGXQNJSKYCXNLZCXM-I20.
Reproducibility Statement
The main paper specifies the model architectures, training setups, and experimental protocols, and the appendix provides additional details on hyperparameters, datasets, preprocessing and data augmentation, training procedures, evaluation methods and scoring rules, robustness and generalization analyses, task specifications, and qualitative results. For full reproducibility, we release:
@@ Acknowledgments
Appendix A LLM Usage
Large Language Models (LLMs) were used only for grammar and wording polishing. All research content and contributions are entirely the responsibility of the authors, with no involvement of LLMs.
Appendix B Robustness and Generalization Evaluation
B.1 Real-world Evaluation
@@ B.2 Simulation Evaluation
Appendix C Additional Training Details
C.1 Hyper-parameters
@@ C.2 Training Data
For the SimplerEnv-Bridge benchmark, we train on BridgeData v2 (walke2023bridgedata), a large-scale, language-conditioned real-robot manipulation dataset of roughly 60,000 teleoperated trajectories collected on WidowX robots across diverse tabletop settings. Episodes pair language instructions with demonstrations of skills such as picking, placing, pushing, stacking, and folding.
@@
For the SimplerEnv-Fractal benchmark, we use RT-1 (brohan2022rt), a large-scale real-world dataset of roughly 130,000 episodes spanning 700+ tasks, collected over 17 months by the Google Robot fleet and paired with natural-language instructions.
@@
LIBERO
LIBERO (liu2023libero) provides simulation tasks with a Franka robot across five suites: Spatial, Object, Goal, Long-10, and Long-90, totaling 130 language-conditioned tasks. Each task supplies 50 demonstrations.
Mikasa-Robo
Mikasa-Robo (cherepanov2025memory) comprises five memory-dependent manipulation tasks, each with 250 officially provided demonstrations, using $\Delta$ end-effector control.
@@ C.3 Training Setup
On Bridge v2, models are trained for 50k steps with a stream dataloader. Each episode is unpacked into consecutive frames tagged with its episode ID. During training, batches are filled sequentially with frames from a single episode whenever possible. If an episode ends before the batch is complete, the remaining slots are filled with frames from the following episode. A new batch then continues from the position where the previous one stopped, ensuring that in-episode temporal order is always preserved. The memory length is fixed to 16.
@@
Following OpenVLA (kim2024openvla), we train with 50 demonstrations per task after removing failed trajectories from the dataset. Spatial, Object, and Goal suites are trained separately for 20k steps each, while Long-10 and Long-90 are treated as a single family of long-horizon data and trained jointly for 40k steps. The dataloader adopts a grouped sampling strategy: in each iteration, 16 frames are randomly sampled from within a single episode, matching the memory length of 16 used throughout training. The dataloader adopts a grouped sampling strategy in which each batch is divided into multiple groups, and each group consists of several frames drawn from a single episode. Frames within a group are kept in temporal order. The memory length is set to 16.
Mikasa-Robo
Following Mikasa-Robo (cherepanov2025memory), we adopt the standard protocol with five tasks and train jointly on all 1,250 demonstrations for 20k steps, using $128\times 128$ RGB observations and $\Delta$ end-effector control. We reuse the same dataloader setup as in LIBERO, and set the memory length to 16.
@@
Models are trained for 5k-20k steps depending on task and dataset size. The general tasks contain 50-150 demonstrations per task, while long-horizon temporal tasks use 200-300 demonstrations per task. The memory length is set to 16 for general tasks and 256 for long-horizon temporal tasks.
C.4 Data Augmentation
@@
Appendix D Additional Evaluation Details
D.1 SimplerEnv
Evaluation follows the official CogACT protocol (li2024cogact). We adopt the same evaluation scripts and use the adaptive action ensemble strategy introduced in CogACT, with ensemble coefficient $\alpha=0.1$, ensemble horizon set to $7$ for Bridge and $2$ for Fractal. For Bridge, models are trained for 50k steps and validated every 2.5k steps, since the denoising objective of diffusion models does not reliably indicate policy quality, we report success rates at the best validation step. For Fractal, training runs for 80k steps with validation every 5k steps, and evaluation covers 336 variants in total (Tab. 19), we similarly report success rates at the best validation step, VM and VA settings are evaluated separately.
@@ D.1 SimplerEnv
D.2 LIBERO
Evaluation on LIBERO (liu2023libero) is conducted across all five suites (Spatial, Object, Goal, Long-10, and Long-90). Models are validated every 1k steps, and each task is evaluated with 50 trials. Success rates are reported at the best validation step. Unlike SimplerEnv, no action ensemble strategy is used in our LIBERO experiments.